OpenStudy (anonymous):

Adding and Subtracting Rational Expressions to Simplest Form (x-5)/5-(x+2)/6+(x+1)/15

OpenStudy (anonymous):

Is this the equation? \[\frac{x-5}{5}-\frac{x+2}{6}+\frac{x+1}{15}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[a*1=a\] \[1 = \frac{1}{1} = \frac{2}{2} = \frac{3}{3} = \frac{4}{4} = \frac{5}{5} = \frac{x}{x}\] Change each number to have the same denominator by multiplying it by 1 in the form of a fraction.